OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $3.95. If the number of dimes was doubled and the number of nickels was increased by 21, the value of the coins would be $8.40. How many nickels and dimes does he have?

OpenStudy (anonymous):

will this one disappear too?

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

\[10x+5y=395\] \[20x+5(y+21)=840\]

OpenStudy (anonymous):

do you know how to solve the system?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not you silly !

OpenStudy (anonymous):

i would rewrite \[20x+5(y+21)=840\] as \[20x+5y=365\] first

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

satellite73 ...step by step explanation ... use elimination ...easier

OpenStudy (anonymous):

then subtract the first equation from the second \[20x+5y=735\] \[10x+5y=395\] subtract to get \[10x=735-395=340\] making \[x=34\]

OpenStudy (anonymous):

is that the anwser?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if you are noot sure ask for more help in the chat

OpenStudy (anonymous):

x=34 means there are 34 dimes. Now, you need to find y to find number of nickels.

OpenStudy (anonymous):

10x+5y=395 x = 34 => 340 + 5y = 395 => 5y = 395-340 => 5y = 55 => y = 11 So, 34 dimes and 11 nickels.

OpenStudy (anonymous):

{5 n + 10 d = 395, 5 (n + 21) + 10 (2 d) = 840} {n = 11, d = 34}